# Date of Acceptance: Nov. 9th 2025

### v1

@@ 1 Introduction
Modern data collection paradigms [DARaiDataset] utilize multifarious sensor modalities to represent an environment, as each modality captures unique aspects of a scene. For example, RGB excels at conveying visual cues in objects such as color or texture, while depth captures geometric structures and directional relationships between objects (Figure 1 (a-b)). Hence, multi-modal learning [yuan2025survey] is indispensable for human-centric tasks such as action anticipation [kim2025multi, zhong2023anticipative, barsoum20203d, guan2022afe, cao2022qmednet, pmlr-v28-koppula13], video understanding [huang2024multi, kaviani2025hierarchical], and action recognition [yarici2025subject], where single modalities fall short in expressing subtle semantic variations within or across scenes. Recent multi-modal fusion methods have achieved empirical success through contrastive losses [radford2021learning, girdhar2023imagebind], attention blocks [yu2019deep], or various specialized objectives [fu2021seamless]. However, they often rely on learning representations via indirect pre-text tasks or alignment losses [liu2025tacfn], rather than directly targeting the informative content in each modality. This lack of targeted information fusion makes multi-modal fusion models susceptible to the common pitfalls: feature collapse  [laurent2023feature] and modality collapse [chaudhuri2025closer]. A central technical challenge we tackle in this paper is the lack of a unifying information framework that targets both feature and modality collapse.
@@
An ideal representation space after fusion should reflect Figure 1(e), where each information component is equally represented, resulting in a space that integrates feature directions encoding useful content from both modalities. However, in reality, different types of collapse can occur in the fused representation space. In Figure 1(f), we visualize feature collapse where certain information components become limited in their contribution to the overall space after fusion. This limits the overall feature diversity and the generalizability of the model [kokilepersaud2025adadim, benkert2024transitional]. In Figure 1(g), we observe modality collapse [chaudhuri2025closer] where information components are expressed for one modality over the other, limiting the sharing of complementary features between the modalities.
To address these challenges, we propose an ideal fusion strategy for action anticipation that addresses both representation collapse. First, to counter feature collapse, we aim to increase the effective rank [roy2007effective] of the fused representation matrix, as it serves as a proxy for the diversity of information. This is motivated by prior work defining the effective rank as matrix entropy [von2018mathematical, renyi1961measures, zhang2023matrix]. Based on this, we develop Rank-enhancing Token fuser, a theoretically grounded fusion framework that selectively blends less informative channels (i.e., channels that contribute little to the principal eigenvector directions) with complementary features from another modality. Second, we posit that compatible modalities mutually increase each other’s effective rank. Hence, to address modality collapse, we perform an analysis of different modalities to pair with RGB for action anticipation. We demonstrate through a relative rank analysis that the depth modality results in the most balanced feature space between both modalities. We use this intuition to develop R3D (Rank-enhancing fusion in 3D), the first depth-informed action anticipation architecture. We demonstrate performance improvements across a wide variety of benchmarks. Our main contributions are summarized as follows:
@@ 2.1 Action Anticipation
The availability of large-scale video datasets [damen2022rescaling, grauman2022ego4d, kuehne2014language, li2018eye, kokilepersaud2023focal] has spurred significant progress in addressing the challenge of action anticipation. These methods can be broadly categorized into two types. First, short-term action anticipation focuses on predicting a single future action that will occur within a few seconds [fernando2021anticipating, furnari2019would, sener2020temporal]. In contrast, long-term action anticipation seeks to forecast an extended sequence of future actions from a long-range video, aiming to predict events far into the future [abu2018will, abu2021long, sener2020temporal, gong2024actfusion, huang2024multi, kim2025multi]. To predict long-term action, recent advances in action anticipation leverage multi-modal inputs to capture complementary contextual cues [lai2024human]. For instance, [zatsarynna2021multi, zhang2024multi] explicitly integrated object features alongside RGB data.  [zhong2023anticipative] leveraged audio to capture off-camera events and to disambiguate visually similar actions.  [beedu2024efficacy, kim2025multi] integrates text to provide rich and fine-grained semantic context. As humans perform actions in a 3D world, prior studies have explored the use of human pose as a complementary modality [duan2022revisiting, guan2022afe, barsoum20203d, cao2022qmednet, pmlr-v28-koppula13], often relying on motion capture systems or pose estimation pipelines. In contrast, we are the first to introduce raw depth data as a multi-modal input without requiring additional motion capture hardware, making it practical and deployable in real-world settings where commodity RGB-D cameras are readily available.
@@ 2.2 Multi-modal Fusion Method
A core challenge in multi-modal learning is the effective fusion of heterogeneous data streams [huang2020pixel, hu2021unit, akkus2023multimodal, chen2024internvl, kim2021vilt]. Fusion strategies are broadly categorized based on their methodology of fusion. Aggregation-based fusion strategies combine features through concatenations, summation, or attention pooling [valada2020self, rotondo2019action, hazirbas2016fusenet, zeng2019deep, prabhushankar2022olives]. Alignment-based fusion methods temporally or semantically align modality representations [wang2016learning, zatsarynna2021multi, zhong2023anticipative, beedu2024efficacy, kim2025multi, huang2020pixel, kim2021vilt]. Finally, hybrid approaches integrate both aggregation and alignment techniques [baltruvsaitis2018multimodal]. Despite their architectural advances, these fusion strategies often suffer from feature redundancy where redundant or weakly informative features from one modality can overwhelm the shared representation space, resulting in suboptimal fusion [liu2025tacfn]. Our method addresses this limitation by encouraging diverse representations, as evidenced by an increased effective rank in the joint embedding space.
@@ 2.3 Approximating Information through Rank
The technical novelty in this paper is targeted multi-modal feature fusion. We utilize the effective rank of the feature matrix as a measure of information content. [von2018mathematical, renyi1961measures] identify that some measure of the uniformity of the singular value spectrum, such as effective rank [roy2007effective], is representative of the entropy of a matrix [zhang2023matrix], and its information content. Intuitively, a more uniform singular value distribution implies that the representation of the data varies along a greater number of feature directions, thus exhibiting greater diversity and countering potential representational collapse [jing2021understanding]. To capture this eigenvalue uniformity as a metric, the effective rank computes the entropy of the normalized eigenspectrum of a matrix $Z$ with associated eigenvalues $\sigma_{1},...,\sigma_{r}]$ as follows:
@@
Due to these useful properties, measuring the rank of a representation space has gained traction in various research areas. Within self-supervised learning, [thilak2023lidar, agrawal2022alpha, garrido2023rankme] use some measure of the eigenvalue spectra as a way to identify better performing models without the need for task specific fine-tuning. [kokilepersaud2025adadim] show that adapting a loss during training based on rank measurements can improve downstream performance. Other works devise loss functions with the explicit goal of increasing rank [bardes2021vicreg, kokilepersaud2024hex] and manipulating eigenvalues directly [kim2023vne]. Despite the potential of integrating rank into training paradigms, there is limited work on applying rank as a mechanism to improve modality fusion. [chaudhuri2025closer] analyze representational collapse in multi-modal architectures through effective rank measurements. However, their method does not involve directly integrating the rank and instead relies on discriminator networks. In contrast, our work is the first to demonstrate that rank information can directly inform targeted feature fusion.
@@ 4.3 Temporal Fuser and Action Anticipation
DARai [kaviani2025hierarchical] (Coarse) Uni RNN [abu2018will] 22.40 22.59 20.71 18.38 30.75 25.34 25.99 23.17
CNN [abu2018will] 13.28 13.70 12.99 13.90 19.85 16.68 18.69 17.48
FUTR [gong2022future] 25.05 25.11 24.48 23.18 40.71 33.57 33.43 30.79
GTAN [zatsarynna2024gated] 27.70 29.13 27.35 26.11 42.47 42.45 42.27 34.37
@@
Multi AFFT [zhong2023anticipative] 23.14 24.78 23.62 21.02 33.82 29.25 28.33 25.45
m&m-Ant [kim2025multi] 25.75 25.70 24.24 23.12 42.00 34.71 34.49 31.34
@@
DARai [kaviani2025hierarchical] (Fine-grained) Uni RNN [abu2018will] 7.45 5.69 4.66 3.58 9.07 7.35 6.40 5.04
CNN [abu2018will] 7.70 6.28 5.38 3.98 6.69 5.48 5.10 3.86
FUTR [gong2022future] 17.09 12.34 10.39 8.18 18.59 15.83 13.26 10.81
GTAN [zatsarynna2024gated] 25.66 24.06 22.60 22.05 26.30 25.43 22.40 18.88
@@
Multi AFFT [zhong2023anticipative] 14.13 13.38 13.49 9.53 17.01 17.26 15.18 10.91
@@
UTKinects [xia2012view] Uni RNN [abu2018will] 24.42 15.37 15.92 11.10 15.22 15.19 13.12 9.89
CNN [abu2018will] 25.00 16.60 14.79 11.07 15.10 16.67 14.29 8.08
FUTR [gong2022future] 29.63 20.75 20.53 16.48 16.46 16.90 14.31 9.95
GTAN [zatsarynna2024gated] 34.38 35.71 39.54 34.17 28.27 32.36 27.22 21.16
@@
Multi AFFT [zhong2023anticipative] 25.00 16.67 16.67 12.50 16.67 16.67 14.29 10.00
@@
NTURGBD [shahroudy2016ntu] Uni RNN [abu2018will] 10.98 10.49 10.53 10.94 8.60 9.55 10.11 11.69
CNN [abu2018will] 15.11 14.91 14.33 15.12 15.09 15.49 15.52 16.74
FUTR [gong2022future] 18.59 18.56 18.61 18.61 20.04 20.13 20.07 20.11
GTAN [zatsarynna2024gated] 18.93 18.51 18.55 18.70 21.63 21.72 21.95 20.93
@@
Multi AFFT [zhong2023anticipative] 20.55 20.76 20.66 20.79 20.72 20.72 20.74 21.27
@@ 5.1 Quantitative Analysis
State-of-the-art comparison. As we observe in Table 1, R3D consistently outperforms the state-of-the-art across all datasets and evaluation settings, achieving an average accuracy gain of up to 3.74%. We compare the performance of multiple architectures for action anticipation and observe that our approach achieves the best performance among all models, with AFFT [zhong2023anticipative], m&m-Ant [kim2025multi], FUTR [gong2022future], and GTAN [zatsarynna2024gated], serving as the state-of-the-art baselines. This gain is particularly notable in scenarios with lower observation rates ($\alpha=0.2$), where depth information allows the model to leverage subtle cues even when visual input is limited. The performance boost is more pronounced on the DARai dataset than on the UTKinects dataset. This is because DARai contains more dense action sequences and more frequent transitions. In this case, depth information enhances the model’s ability to capture nuanced temporal variations, leading to greater improvements compared to other datasets. In contrast, on NTURGBD, the strength of R3D is less pronounced. This can be attributed to the fact that NTURGBD provides high-quality RGB-Depth pairs with well-trimmed and well-curated action segments. While R3D is designed to be robust in less curated, noisy real-world settings (See 5.4), its relative advantage diminishes when both modalities are already strong, which is less common in real-world settings.
@@ 5.3 More SOTA comparisons
RGB - Depth AFFT [zhong2023anticipative] 14.13 13.38 13.49 9.53 17.01 17.26 15.18 10.91
R3D (Ours) 30.57 25.92 23.02 16.68 26.97 26.43 23.81 18.02
Different-view RGBs AFFT [zhong2023anticipative] 11.53 8.46 7.47 5.84 10.61 8.51 8.22 6.88
@@
RGB - Text AFFT [zhong2023anticipative] 20.21 21.83 21.93 17.07 25.79 25.62 24.90 19.62
@@
More modality pairs. To further validate the general effectiveness of R3D with different modalities, as shown in Table 4, we additionally present the State-of-the-art comparisons of action anticipation performance across four different modality pairs: Depth, Multi-view RGB, IMU, and Text at varying observation rates ($\alpha$) and prediction rates ($\beta$). For comparison, we use AFFT [zhong2023anticipative], the current state-of-the-art method in multi-modal action anticipation. To ensure fairness and reproducibility, we conduct three runs with fixed random seeds (1, 10, 13452) and simply replace the input from the Figure 4. Overall, R3D achieves superior performance to AFFT for modality pairs with higher harmonic means as illustrated in Figure 3. For RGB - IMU pair, R3D underperforms compared to AFFT. We attribute this to the lack of sufficiently balanced interaction between RGB and IMU features, as also reflected in Figure 3. We expect that tailoring the fusion parameters for RGB - IMU would enhance the interaction and further improve performance.
@@
DARai [kaviani2025hierarchical] (Coarse) ActFusion [gong2024actfusion] 31.38 26.62
@@
DARai [kaviani2025hierarchical] (Fine-grained) ActFusion [gong2024actfusion] 17.83 18.16
@@
NTURGBD [shahroudy2016ntu] ActFusion [gong2024actfusion] 13.26 6.00
@@
More downstream tasks. To further validate the general effectiveness of RTF, we additionally apply our method to Action Segmentation task. Unlike Action Anticipation, Action Segmentation task segments an input video into temporally contiguous segments, assigning a categorical action label to each segment independently [lea2017temporal, ishikawa2021alleviating, farha2019ms, wang2020boundary, yi2021asformer, xu2022don, behrmann2022unified, liu2023diffusion, gong2024actfusion]. To comprehensively evaluate the task, first, we assess the correctness of predictions at every individual frame by reporting frame-wise accuracy [lea2017temporal, ishikawa2021alleviating, wang2020boundary, xu2022don, behrmann2022unified, gong2024actfusion]. Additionally, we also measure how well the predicted segments align with the true temporal boundaries and labels of actions by measuring an edit score, a segment-wise metrics [farha2019ms, yi2021asformer, liu2023diffusion, gong2024actfusion]. To obtain results from the action segmentation task, we use the output of the Segmentation Head of the Temporal Fuser of R3D, illustrated in Figure 4 and explained in detail in Section C.1. To ensure consistency, we report the average performance over three runs with fixed random seeds (1, 10, 13452).

### v2

@@ 1 Introduction
Modern data collection paradigms [33] utilize multifarious sensor modalities to represent an environment, as each modality captures unique aspects of a scene. For example, RGB excels at conveying visual cues in objects such as color or texture, while depth captures geometric structures and directional relationships between objects (Figure 1 (a-b)). Hence, multi-modal learning [70] is indispensable for human-centric tasks such as action anticipation [36, 76, 7, 25, 11, 41], video understanding [28, 32], and action recognition [67], where single modalities fall short in expressing subtle semantic variations within or across scenes. Recent multi-modal fusion methods have achieved empirical success through contrastive losses [54, 21], attention blocks [69], or various specialized objectives [18]. However, they often rely on learning representations via indirect pre-text tasks or alignment losses [48], rather than directly targeting the informative content in each modality. This lack of targeted information fusion makes multi-modal fusion models susceptible to the common pitfalls: feature collapse  [44] and modality collapse [12]. A central technical challenge we tackle in this paper is the lack of a unifying information framework that targets both feature and modality collapse.
@@
An ideal representation space after fusion should reflect Figure 1(e), where each information component is equally represented, resulting in a space that integrates feature directions encoding useful content from both modalities. However, in reality, different types of collapse can occur in the fused representation space. In Figure 1(f), we visualize feature collapse where certain information components become limited in their contribution to the overall space after fusion. This limits the overall feature diversity and the generalizability of the model [40, 10]. In Figure 1(g), we observe modality collapse [12] where information components are expressed for one modality over the other, limiting the sharing of complementary features between the modalities.
To address these challenges, we propose an ideal fusion strategy for action anticipation that addresses both representation collapse. First, to counter feature collapse, we aim to increase the effective rank [57] of the fused representation matrix, as it serves as a proxy for the diversity of information. This is motivated by prior work defining the effective rank as matrix entropy [62, 55, 74]. Based on this, we develop Rank-enhancing Token fuser, a theoretically grounded fusion framework that selectively blends less informative channels (i.e., channels that contribute little to the principal eigenvector directions) with complementary features from another modality. Second, we posit that compatible modalities mutually increase each other’s effective rank. Hence, to address modality collapse, we perform an analysis of different modalities to pair with RGB for action anticipation. We demonstrate through a relative rank analysis that the depth modality results in the most balanced feature space between both modalities. We use this intuition to develop R3D (Rank-enhancing fusion in 3D), the first depth-informed action anticipation architecture. We demonstrate performance improvements across a wide variety of benchmarks. Our main contributions are summarized as follows:
@@ 2.1 Action Anticipation
The availability of large-scale video datasets [14, 24, 42, 46, 38] has spurred significant progress in addressing the challenge of action anticipation. These methods can be broadly categorized into two types. First, short-term action anticipation focuses on predicting a single future action that will occur within a few seconds [17, 19, 58]. In contrast, long-term action anticipation seeks to forecast an extended sequence of future actions from a long-range video, aiming to predict events far into the future [1, 2, 58, 23, 28, 36]. To predict long-term action, recent advances in action anticipation leverage multi-modal inputs to capture complementary contextual cues [43]. For instance, [71, 75] explicitly integrated object features alongside RGB data.  [76] leveraged audio to capture off-camera events and to disambiguate visually similar actions.  [8, 36] integrates text to provide rich and fine-grained semantic context. As humans perform actions in a 3D world, prior studies have explored the use of human pose as a complementary modality [15, 25, 7, 11, 41], often relying on motion capture systems or pose estimation pipelines. In contrast, we are the first to introduce raw depth data as a multi-modal input without requiring additional motion capture hardware, making it practical and deployable in real-world settings where commodity RGB-D cameras are readily available.
@@ 2.2 Multi-modal Fusion Method
A core challenge in multi-modal learning is the effective fusion of heterogeneous data streams [29, 27, 4, 13, 37]. Fusion strategies are broadly categorized based on their methodology of fusion. Aggregation-based fusion strategies combine features through concatenations, summation, or attention pooling [61, 56, 26, 73, 53]. Alignment-based fusion methods temporally or semantically align modality representations [63, 71, 76, 8, 36, 29, 37]. Finally, hybrid approaches integrate both aggregation and alignment techniques [5]. Despite their architectural advances, these fusion strategies often suffer from feature redundancy where redundant or weakly informative features from one modality can overwhelm the shared representation space, resulting in suboptimal fusion [48]. Our method addresses this limitation by encouraging diverse representations, as evidenced by an increased effective rank in the joint embedding space.
@@ 2.3 Approximating Information through Rank
The technical novelty in this paper is targeted multi-modal feature fusion. We utilize the effective rank of the feature matrix as a measure of information content. [62, 55] identify that some measure of the uniformity of the singular value spectrum, such as effective rank [57], is representative of the entropy of a matrix [74], and its information content. Intuitively, a more uniform singular value distribution implies that the representation of the data varies along a greater number of feature directions, thus exhibiting greater diversity and countering potential representational collapse [31]. To capture this eigenvalue uniformity as a metric, the effective rank computes the entropy of the normalized eigenspectrum of a matrix $Z$ with associated eigenvalues $\sigma_{1},...,\sigma_{r}]$ as follows:
@@
Due to these useful properties, measuring the rank of a representation space has gained traction in various research areas. Within self-supervised learning, [60, 3, 20] use some measure of the eigenvalue spectra as a way to identify better performing models without the need for task specific fine-tuning. [40] show that adapting a loss during training based on rank measurements can improve downstream performance. Other works devise loss functions with the explicit goal of increasing rank [6, 39] and manipulating eigenvalues directly [35]. Despite the potential of integrating rank into training paradigms, there is limited work on applying rank as a mechanism to improve modality fusion. [12] analyze representational collapse in multi-modal architectures through effective rank measurements. However, their method does not involve directly integrating the rank and instead relies on discriminator networks. In contrast, our work is the first to demonstrate that rank information can directly inform targeted feature fusion.
@@ 4.3 Temporal Fuser and Action Anticipation
DARai [32] (Coarse) Uni RNN [1] 22.40 22.59 20.71 18.38 30.75 25.34 25.99 23.17
CNN [1] 13.28 13.70 12.99 13.90 19.85 16.68 18.69 17.48
FUTR [22] 25.05 25.11 24.48 23.18 40.71 33.57 33.43 30.79
GTAN [72] 27.70 29.13 27.35 26.11 42.47 42.45 42.27 34.37
@@
Multi AFFT [76] 23.14 24.78 23.62 21.02 33.82 29.25 28.33 25.45
m&m-Ant [36] 25.75 25.70 24.24 23.12 42.00 34.71 34.49 31.34
@@
DARai [32] (Fine-grained) Uni RNN [1] 7.45 5.69 4.66 3.58 9.07 7.35 6.40 5.04
CNN [1] 7.70 6.28 5.38 3.98 6.69 5.48 5.10 3.86
FUTR [22] 17.09 12.34 10.39 8.18 18.59 15.83 13.26 10.81
GTAN [72] 25.66 24.06 22.60 22.05 26.30 25.43 22.40 18.88
@@
Multi AFFT [76] 14.13 13.38 13.49 9.53 17.01 17.26 15.18 10.91
@@
UTKinects [65] Uni RNN [1] 24.42 15.37 15.92 11.10 15.22 15.19 13.12 9.89
CNN [1] 25.00 16.60 14.79 11.07 15.10 16.67 14.29 8.08
FUTR [22] 29.63 20.75 20.53 16.48 16.46 16.90 14.31 9.95
GTAN [72] 34.38 35.71 39.54 34.17 28.27 32.36 27.22 21.16
@@
Multi AFFT [76] 25.00 16.67 16.67 12.50 16.67 16.67 14.29 10.00
@@
NTURGBD [59] Uni RNN [1] 10.98 10.49 10.53 10.94 8.60 9.55 10.11 11.69
CNN [1] 15.11 14.91 14.33 15.12 15.09 15.49 15.52 16.74
FUTR [22] 18.59 18.56 18.61 18.61 20.04 20.13 20.07 20.11
GTAN [72] 18.93 18.51 18.55 18.70 21.63 21.72 21.95 20.93
@@
Multi AFFT [76] 20.55 20.76 20.66 20.79 20.72 20.72 20.74 21.27
@@ 5.1 Quantitative Analysis
State-of-the-art comparison. As we observe in Table 1, R3D consistently outperforms the state-of-the-art across all datasets and evaluation settings, achieving an average accuracy gain of up to 3.74%. We compare the performance of multiple architectures for action anticipation and observe that our approach achieves the best performance among all models, with AFFT [76], m&m-Ant [36], FUTR [22], and GTAN [72], serving as the state-of-the-art baselines. This gain is particularly notable in scenarios with lower observation rates ($\alpha=0.2$), where depth information allows the model to leverage subtle cues even when visual input is limited. The performance boost is more pronounced on the DARai dataset than on the UTKinects dataset. This is because DARai contains more dense action sequences and more frequent transitions. In this case, depth information enhances the model’s ability to capture nuanced temporal variations, leading to greater improvements compared to other datasets. In contrast, on NTURGBD, the strength of R3D is less pronounced. This can be attributed to the fact that NTURGBD provides high-quality RGB-Depth pairs with well-trimmed and well-curated action segments. While R3D is designed to be robust in less curated, noisy real-world settings (See 5.4), its relative advantage diminishes when both modalities are already strong, which is less common in real-world settings.
@@ 5.3 More SOTA comparisons
RGB - Depth AFFT [76] 14.13 13.38 13.49 9.53 17.01 17.26 15.18 10.91
R3D (Ours) 32.57 25.92 24.02 16.68 26.97 28.43 24.81 18.02
Different-view RGBs AFFT [76] 11.53 8.46 7.47 5.84 10.61 8.51 8.22 6.88
@@
RGB - Text AFFT [76] 20.21 21.83 21.93 17.07 25.79 25.62 24.90 19.62
@@
More modality pairs. To further validate the general effectiveness of R3D with different modalities, as shown in Table 4, we additionally present the State-of-the-art comparisons of action anticipation performance across four different modality pairs: Depth, Multi-view RGB, IMU, and Text at varying observation rates ($\alpha$) and prediction rates ($\beta$). For comparison, we use AFFT [76], the current state-of-the-art method in multi-modal action anticipation. To ensure fairness and reproducibility, we conduct three runs with fixed random seeds (1, 10, 13452) and simply replace the input from the Figure 4. Overall, R3D achieves superior performance to AFFT for modality pairs with higher harmonic means as illustrated in Figure 3. For RGB - IMU pair, R3D underperforms compared to AFFT. We attribute this to the lack of sufficiently balanced interaction between RGB and IMU features, as also reflected in Figure 3. We expect that tailoring the fusion parameters for RGB - IMU would enhance the interaction and further improve performance.
@@
DARai [32] (Coarse) ActFusion [23] 31.38 26.62
@@
DARai [32] (Fine-grained) ActFusion [23] 17.83 18.16
@@
NTURGBD [59] ActFusion [23] 13.26 6.00
@@
More downstream tasks. To further validate the general effectiveness of RTF, we additionally apply our method to Action Segmentation task. Unlike Action Anticipation, Action Segmentation task segments an input video into temporally contiguous segments, assigning a categorical action label to each segment independently [45, 30, 16, 64, 68, 66, 9, 47, 23]. To comprehensively evaluate the task, first, we assess the correctness of predictions at every individual frame by reporting frame-wise accuracy [45, 30, 64, 66, 9, 23]. Additionally, we also measure how well the predicted segments align with the true temporal boundaries and labels of actions by measuring an edit score, a segment-wise metrics [16, 68, 47, 23]. To obtain results from the action segmentation task, we use the output of the Segmentation Head of the Temporal Fuser of R3D, illustrated in Figure 4 and explained in detail in Section C.1. To ensure consistency, we report the average performance over three runs with fixed random seeds (1, 10, 13452).